65,98,55,62,79,59,51,90,72,56,70,62,66,80,94,79,63,73,71,85


Organize the number of oil changes into a frequency distribution


1
Expert's answer
2021-06-15T06:30:14-0400
"51,55,56,59,62,62,63,65,66,70,"

"71,72,73,79,79,80,85,90,94,98"

Sample size: "n=20."

Maximum value "Max= 98."

Minimum value "Min= 51."

Measurement "MU=1"


Amount of classes


"k=1+3.322\log_{10}(n)"

"=1+3.322\log_{10}(20)\approx5.322"

We must approach to the upper integer: "k=6."

Class interval


"CI=\dfrac{98-51+1}{6}=8"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Class & Class & Frequency \\
interval & boundary & f\\ \hline
 51-58 & 50.5-58.5 & 3 \\
 \hdashline
 59-66 & 58.5-66.5 & 6 \\
 \hdashline
 67-74 & 66.5-74.5 & 4 \\
 \hdashline
 75-82 & 74.5-82.5 & 3 \\
 \hdashline
 83-90 & 82.5-90.5 & 2 \\
 \hdashline
 91-98 & 90.5-98.5 & 2 \\
 
 \hline
\end{array}"